OpenStudy (anonymous):

Solve: 25x2 − 36 = 0.

OpenStudy (ali2x2):

1 x = 6/5 = 1.200 2 x = -6/5 = -1.200

OpenStudy (ali2x2):

2 solutions

OpenStudy (anonymous):

rocker